OpenStudy (anonymous):

Solve algebraically (without using logarithms): (1/8)^2x = 16^(x-5)

OpenStudy (anonymous):

The way to solve it without logarithms is to have the same base.

TheSmartOne (thesmartone):

Hi there! @lisamath12 Nice to meet you again! :)

TheSmartOne (thesmartone):

Ok, so we need the same base, because we cannot solve this using logarithms. And the only other way it to get the same base. Do you agree? :)

TheSmartOne (thesmartone):

Ok, I'll give you a hint since we don't want you to be stuck at the same spot for forever. :) \(\sf\Large \frac{1}{8} = \frac{1}{2^3} = 2^{-3}\) Second hint: \(\sf\Large 16 = 2^4\) Can you use this information to make both sides of the equation have the same base? Remember, if \(\sf a^b = a^c\) then that means b = c

OpenStudy (anonymous):

@lisamath12 Did you get anything that I have said above? :)

TheSmartOne (thesmartone):

@lisamath12 Hello? No response from you, and it's been 2 days.